A large truck is carrying a heavy load. A significant hazard for the driver is that the load may slide forward, crushing the cab, if the truck stops suddenly in an accident or even in braking. Assume a 10,000 kg load sits on the flatbed of a 20,000 kg truck moving at 12 m/s. Assume the load is not tied down to the truck but has a coefficient of friction 0.5 with the flatbed of the truck. Calculate the minimum stopping distance for which the load will not slide forward relative to the truck.
Answer #1

Solution)

Here, Force exerted by the truck= Frictional force

So, ma= uFn= umg

Hence, a= ug=0.5*9.8= 4.9 m/s^2

Now, we know

V^2=u^2+ 2as

0=12^2+2(-4.9)*s

Hence, distance s= 14.69 m

=========